Common stock value—Variable growth Newman Manufacturing is considering a cash purchase of the stock of Grips Tool. During the(Round to the nearest cent)
Answer #1

D0 = $1.92

Growth rate for next 3 years is 30% and a constant growth rate (g) of 9% thereafter

D1 = $1.92 * 1.30 = $2.496
D2 = $2.496 * 1.30 = $3.2448
D3 = $3.2448 * 1.30 = $4.21824
D4 = $4.21824 * 1.09 = $4.5978816

Required Return, rs = 14%

P3 = D4 / (rs - g)
P3 = $4.5978816 / (0.14 - 0.09)
P3 = $4.5978816 / 0.05
P3 = $91.957632

P0 = $2.496/1.14 + $3.2448/1.14^2 + $4.21824/1.14^3 + $91.957632/1.14^3
P0 = $69.60

So, Newman Manufacturing should pay a maximum price of $69.60 per share.